OpenStudy (anonymous):

what is the answer to (3-x)(5-x) divided by (x-3)(x-5)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

ans is 1

OpenStudy (anonymous):

(3-x)(5-x)=-(x-3)*-(x-5)=(x-3)(x-5) 2 minus=+ so u divide something by itself, implies answer is 1

OpenStudy (anonymous):

Yes Andras has given the correct solution

OpenStudy (anonymous):

michael4001, pls close the question by clicking on GOOD ANSWER button